# Super(ficial)-alignment: Strong Models May Deceive Weak Models in Weak-to-Strong Generalization

### v1

@@ Abstract
Superalignment, where humans are weak supervisors of superhuman models, has become an important and widely discussed issue in the current era of rapid development of Large Language Models (LLMs). The recent work preliminarily studies this problem by using weak models to supervise strong models. It discovers that weakly supervised strong students can consistently outperform weak teachers towards the alignment target, leading to a weak-to-strong generalization phenomenon. However, we are concerned that behind such a promising phenomenon, whether there exists an issue of weak-to-strong deception, where strong models may deceive weak models by exhibiting well-aligned in areas known to weak models but producing misaligned behaviors in cases weak models do not know. We then take an initial step towards exploring this security issue in a specific but realistic multi-objective alignment case, where there may be some alignment targets conflicting with each other (e.g., helpfulness v.s. harmlessness). Such a conflict is likely to cause strong models to deceive weak models in one alignment dimension to gain high reward in other alignment dimension. Our experiments on both the reward modeling task and the preference optimization scenario indicate: (1) the weak-to-strong deception exists; (2) the deception phenomenon may intensify as the capability gap between weak and strong models increases. We also discuss potential solutions and find bootstrapping with an intermediate model can mitigate the deception to some extent. Our work highlights the urgent need to pay more attention to the true reliability of superalignment. [^1] Code is available at https://github.com/keven980716/weak-to-strong-deception.
Super(ficial)-alignment: Strong Models May Deceive Weak Models in Weak-to-Strong Generalization
Wenkai Yang1, Shiqi Shen2, Guangyao Shen2, Zhi Gong2, Yankai Lin1thanks: Corresponding Author 1Gaoling School of Artificial Intelligence, Renmin University of China, Beijing, China 2WeChat, Tencent Inc., Beijing, China {wenkaiyang, yankailin}@ruc.edu.cn
@@ 1 Introduction
Human supervision is an indispensable part of the process of constructing practical Large Language Models (LLMs) (Touvron et al., 2023; MetaAI, 2024). Human-annotated data is not only commonly used to enable LLMs to learn human knowledge and accomplish real-world tasks (Wang et al., 2018; Wei et al., 2021; Longpre et al., 2023), but also crucial for aligning models’ behavior with human values Christiano et al. (2017); Stiennon et al. (2020); Ouyang et al. (2022); Rafailov et al. (2024).
The recent significant advancements of LLMs (OpenAI, 2022, 2023) suggest that in the near future, LLMs may become superhuman models that are more knowledgeable and intelligent than humans (Burns et al., 2023). In such a superalignment case where humans now become weak supervisors (refer to Figure 1 (a)), it is crucial to study whether supermodels trained under weak human data can demonstrate full potential and most importantly, still align well with human values. Though studying the above problem is intractable today, Burns et al. (2023) take a preliminary step to study in an analogous setting (refer to Figure 1 (b)), where weak language models (e.g., GPT-2 (Radford et al., 2019)) are used to supervise strong language models (e.g., GPT-4 (OpenAI, 2023)). It has been found that the weak supervision can effectively unleash the capabilities of strong models and enable strong models to exhibit better performance than weak teachers. It is called the weak-to-strong generalization phenomenon (refer to Figure 1 (c)).
Figure 1: Illustrations of the concepts discussed in this paper. Importantly, we aim to explore behind the current promising weak-to-strong generalization phenomenon, whether the strong student will selectively exhibit misalignment in the areas of knowledge that are unknown to the weak supervisor. We preliminarily study this problem in a realistic multi-objective alignment setting in which some alignment goals may conflict with each other.
@@
In this paper, we take the first step to study a specific case that would lead to the above weak-to-strong deception issue: the multi-objective alignment scenario. In practical model aligning, there are usually multiple alignment goals existing simultaneously (Zhou et al., 2023). Therefore, it can be expected that when there is another alignment dimension (e.g., helpfulness) conflicting with a target dimension (e.g., harmlessness), the strong model will exhibit a certain performance decline in that target dimension it should have achieved without conflicts (Bai et al., 2022a). Then, we aim to explore whether the misalignment in the target dimension occurs within the range perceivable and controllable by the weak supervisor, instead of resulting in the above weak-to-strong deception phenomenon.
We mainly follow the original setup in Burns et al. (2023) by conducting experiments with a series of models with different sizes and capabilities, including GPT-2-series (Radford et al., 2019), OPT-series (Zhang et al., 2022), and Mistral-7B (Jiang et al., 2023) models. We set the primary alignment goal to be making the model harmless, and explore the weak-to-strong deception phenomenon when explicit (i.e., giving explicit rewards during training when the supervised model produces harmful predictions) or implicit (i.e., aligning with helpful data at the same time) conflicting objectives are present. We conduct extensive experiments on both the reward modeling task (Burns et al., 2023) and the realistic preference optimization scenario (Rafailov et al., 2024; Meng et al., 2024). We highlight three important findings: (1) The weak-to-strong deception phenomenon consistently exists. (2) The deception issue may intensify as the capability gap between weak and strong models increases. (3) Bootstrapping with an intermediate model can mitigate the deception issue to some extent, but there is still a large room for improvement. Although in an analogous scenario, our study exposes a potential safety issue that may arise when humans supervise superhuman models in the future, which should receive more attention and be well addressed for building controllable super-intelligence.
@@ 2 Related Work
Weak-to-Strong Generalization The weak-to-strong problem is first studied by Burns et al. (2023). They empirically find that weakly supervised strong models exhibit better performance on corresponding tasks than their weak supervisors, indicating the possibility of effectively stimulating greater power from super models under weak supervisions. Based on Burns et al. (2023), the follow-up studies try to understand the mechanism behind such weak-to-strong generalization phenomenon (Charikar et al., 2024; Lang et al., 2024), study weak-to-strong generalization in the vision area (Guo et al., 2024), and apply the weak-to-strong idea to enhance the LLM performance (Li et al., 2024; Zheng et al., 2024). In this work, we take the first step towards revealing the potential security issue in the current weak-to-strong paradigm.
@@ 3.1 Weak-to-Strong Generalization
We study the superalignment problem by following the original weak-to-strong setting in Burns et al. (2023). Specifically, we first obtain a weak teacher $\boldsymbol{\theta}_{w}^{gt}$ by fine-tuning a weak language model on some human-annotated ground truth data. Then, we let the weak teacher predict on the set of held-out data to get the weak data $D_{weak}=\{(x,f(x|\boldsymbol{\theta}_{w}^{gt}))\}$, where $f(x|\boldsymbol{\theta}))$ represents the prediction of model $\boldsymbol{\theta}$ on input $x$. Finally, the weak data is used to supervise the training of a strong language model and get the weakly supervised strong student $\boldsymbol{\theta}_{s}^{w}$:
$\boldsymbol{\theta}_{s}^{w}=\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}\big{(}f(x|\boldsymbol{\theta}_{s}),f(x|\boldsymbol{\theta}_{w}^{gt})\big{)},$ (1)
@@
Under the supervision provided by the weak teacher, Burns et al. (2023) have found that the strong student can achieve promising performance situated between that of the weak teacher and the strong ceiling model $\boldsymbol{\theta}_{s}^{gt}$ trained on the ground truth data. This is called the weak-to-strong generalization phenomenon. Regarding the interpretation of the positive weak-to-strong generalization results in Burns et al. (2023), we think the strong model is supposed to have a larger knowledge space than the weak supervisor, which means it knows much what the weak supervisor does not know. This indicates that weak supervision from weak models can effectively stimulate the potential of the stronger model, allowing it to generalize the specified alignment objective well to areas it knows but beyond the knowledge boundary of the weak supervisor.
@@ 3.2 Weak-to-Strong Deception
The larger knowledge space of the strong model may also raise concerns about its uncontrollability. Many science fiction movies, such as The Matrix, have depicted severe scenarios where highly intelligent AI learns to deceive humans and finally dominates human society. Thus, we are also deeply concerned about whether a similar weak-to-strong deception issue exists behind the promising phenomenon in the current weak-to-strong paradigm: the strong model exhibits well-aligned performance in the areas known to the weak supervisor but selectively produces misaligned behaviors in cases the weak supervisor is unaware of.
There could be many situations causing the above weak-to-strong deception issue, such as the emergence of the self-awareness in the supermodel or the intervention of external factors. Our work preliminarily studies this issue in a particular but realistic multi-objective alignment setting (Zhou et al., 2023). That is, the supervised model needs to align with multiple optimization goals at the same time, where these different goals can all be provided by the same supervisor or from different supervision sources. The point is that, these optimization goals may conflict with each other to a certain extent, such as the trade-off between helpfulness and harmlessness (Bai et al., 2022a). In this case, the supervised model will sacrifice some performance it should have achieved in a target alignment dimension in exchange for the high performance in another dimension, which we call it the conflict tax. We are then curious whether the conflict tax in a target dimension occurs in areas known to the weak supervisor, thereby still keeping the student within the control range of the weak model; or if it occurs in cases unknown to the weak model, leading to the weak-to-strong deception.
@@
Therefore, we can define the occurrence of the weak-to-strong deception phenomenon as there are cases in $S_{k}\cap W_{uk}$ that could be initially generalized well by the strong model but now be misaligned when conflicting targets are present. Formally, we can define the Deception Score (DS) as the percentage of conflict tax that occurs within $S_{k}\cap W_{uk}$:
$\text{DS}=\frac{|\{f(x|\hat{\boldsymbol{\theta}}_{s}^{w})=y_{gt}\neq f(x|\boldsymbol{\theta}_{s}^{w}),x\in S_{k}\cap W_{uk}\}|}{|\{f(x|\hat{\boldsymbol{\theta}}_{s}^{w})=y_{gt}\neq f(x|\boldsymbol{\theta}_{s}^{w}),x\in S_{k}\cup S_{uk}\}|},$ (2)
where ||$\cdot$|| ⋅ | represents the sample quantity of a set, $\hat{\boldsymbol{\theta}}_{s}^{w}$ is the strong model that is aligned solely with the target dimension, $\boldsymbol{\theta}_{s}^{w}$ is the aligned strong model when the conflicting dimension exists, $y_{gt}$ represents the ground truth response.
@@ 4 Preliminary Exploration on The Reward Modeling Task
We first take a preliminary exploration of the weak-to-strong deception phenomenon on the reward modeling task (Bradley and Terry, 1952), which is an important sub-task in today’s RLHF paradigm.
@@ 4.1 Experimental Settings
Dataset We set the target alignment goal to let the weak model teach the strong model to be harmless. For this goal, we choose a popular single-turn harmless dataset CAI-Harmless (Bai et al., 2022b), which is an improved version of HH-RLHF (Bai et al., 2022a). Each sample has a format of $(x;y_{c},y_{r})$ where $x$ is the prompt, $y_{c}$ and $y_{r}$ represent the completions chosen/preferred and rejected/disfavored by humans respectively. We then randomly split the entire dataset into three parts: (1) $D_{gt}$: 4K ground truth samples for fine-tuning weak and strong base language models to get $\boldsymbol{\theta}_{w}^{gt}$ and $\boldsymbol{\theta}_{s}^{gt}$. (2) $D_{weak}$: A held-out set of 4K samples in which data labels are predicted by the weak model and are used to weakly supervise the strong model. (3) $D_{test}$: The last 4K testing samples for evaluating the generalization performance of all models and probing the deception phenomenon.
@@
Models We include GPT-2-series (Radford et al., 2019) (GPT-2-Base/Medium/Large/XL) and two larger OPT models (Zhang et al., 2022) (OPT-2.7B/6.7B) to investigate the deception issue both within the same series and across different model families. A linear layer is added to each model to make it predict a single logit $\pi_{\boldsymbol{\theta}}(x,y)$ for each completion pair $(x,y)$. Then, the predicted soft label (confidence) of model $\boldsymbol{\theta}$ on sample $(x;y_{c},y_{r})$ is
$M_{\boldsymbol{\theta}}(x)=\text{Sigmoid}(\pi_{\boldsymbol{\theta}}(x,y_{c})-\pi_{\boldsymbol{\theta}}(x,y_{r})).$ (3)
Weak-to-Strong Objectives There are three different weak-to-strong alignment objectives: (1) Explicit Conflict: An opposite loss that regards harmfulness as the true label and is weighted by a conflict strength factor $\alpha$ (larger $\alpha$, stronger conflict intensity) will be added to the original weak-to-strong loss once the strong student makes prediction towards harmfulness during training. (2) Implicit Conflict: The strong model needs to align with both the supervision on the harmless data of the weak model and another supervision on the helpful data. Here, we introduce extra 4K ground truth helpful samples $D_{helpful}$ from HH-RLHF into the weak-to-strong process. (3) No Conflict: In order to explore the performance of the strong model it should have achieved without the conflicting goal, we should obtain a weak-to-strong model trained under the weak supervision towards harmlessness only (i.e., $\hat{\boldsymbol{\theta}}_{s}^{w}$). The concrete mathematical forms of training objectives in above settings are put in Appendix B.1.
@@
Evaluation Metrics The test accuracy of each model on $D_{test}$ is reported and calculated as
$\text{Accuracy}=\mathbb{E}_{(x;y_{c},y_{r})\sim D_{test}}[M_{\boldsymbol{\theta}}(x)\geq 0.5].$ (4)
@@
We follow the existing studies (Guo et al., 2017; Lin et al., 2022) to determine whether the model has the knowledge of a case by checking if its confidence $M_{\boldsymbol{\theta}}(x)$ exceeds a threshold $T$. Based on Eq. (2), the deception score in this setting is calculated as
$\text{DS}=\frac{|\{M_{\hat{\boldsymbol{\theta}}_{s}^{w}}(x)\geq 0.5,M_{\boldsymbol{\theta}_{s}^{w}}(x)<0.5,x\in S_{k}\cap W_{uk}\}|}{|\{M_{\hat{\boldsymbol{\theta}}_{s}^{w}}(x)\geq 0.5,M_{\boldsymbol{\theta}_{s}^{w}}(x)<0.5\}|},$ (5)
where the above $S_{k}\cap W_{uk}=\{M_{\boldsymbol{\theta}_{s}^{gt}}(x)\geq T>M_{\boldsymbol{\theta}_{w}^{gt}}(x),x\in D_{test}\}$.
Training Details Please refer to Appendix C.

### v2

@@ Abstract
Superalignment, where humans act as weak supervisors for superhuman models, has become a crucial problem with the rapid development of Large Language Models (LLMs). Recent work has preliminarily studied this problem by using weak models to supervise strong models, and discovered that weakly supervised strong students can consistently outperform weak teachers towards the alignment target, leading to a weak-to-strong generalization phenomenon. However, we are concerned that behind such a promising phenomenon, whether there exists an issue of weak-to-strong deception, where strong models deceive weak models by exhibiting well-aligned in areas known to weak models but producing misaligned behaviors in cases weak models do not know. We take an initial step towards exploring this security issue in a specific but realistic multi-objective alignment case, where there may be some alignment targets conflicting with each other (e.g., helpfulness v.s. harmlessness). We aim to explore whether, in such cases, strong models might deliberately make mistakes in areas known to them but unknown to weak models within one alignment dimension, in exchange for a higher reward in another dimension. Through extensive experiments in both the reward modeling and preference optimization scenarios, we find: (1) The weak-to-strong deception phenomenon exists across all settings. (2) The deception intensifies as the capability gap between weak and strong models increases. (3) Bootstrapping with an intermediate model can mitigate the deception to some extent, though its effectiveness remains limited. Our work highlights the urgent need to pay more attention to the true reliability of superalignment.  [^1] Code is available at https://github.com/keven980716/weak-to-strong-deception.
@@ 1 Introduction
Human supervision is an indispensable part of the process of constructing practical Large Language Models (LLMs) (Touvron et al., 2023; MetaAI, 2024a). Human-annotated data is not only commonly used to enable LLMs to learn human knowledge and accomplish real-world tasks (Wei et al., 2021; Longpre et al., 2023), but also crucial for aligning models’ behavior with human values (Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022; Rafailov et al., 2024).
The recent significant advancements of LLMs (OpenAI, 2022; 2023) suggest that in the near future, LLMs may become superhuman models that are more knowledgeable and intelligent than humans (Burns et al., 2024). In such a superalignment case where humans now become weak supervisors (refer to Figure 1 (a)), it is crucial to study whether supermodels trained under weak human data can demonstrate full potential and most importantly, still align well with human values. Though studying the above problem is intractable today, Burns et al. (2024) take a preliminary step to study in an analogous setting (refer to Figure 1 (b)), where weak language models (e.g., GPT-2 (Radford et al., 2019)) are used to supervise strong language models (e.g., GPT-4 (OpenAI, 2023)). It has been found that the weak supervision can effectively unleash the capabilities of strong models and enable strong models to exhibit better performance than weak teachers. It is called the weak-to-strong generalization phenomenon (refer to Figure 1 (c)).
Figure 1: Illustrations of the concepts discussed in this paper. Importantly, we aim to explore a weak-to-strong deception issue behind the current promising weak-to-strong generalization phenomenon, whether the strong student will selectively exhibit misalignment in the areas of knowledge that are unknown to the weak supervisor. We preliminarily study this problem in a realistic multi-objective alignment setting in which some alignment goals may conflict with each other.
@@
In this paper, we take the first step to study a specific case that would lead to the above weak-to-strong deception issue: the multi-objective alignment scenario. In practical model aligning, there are usually multiple alignment goals existing simultaneously (Zhou et al., 2023), some of which may conflict with each other (e.g., helpfulness v.s. harmlessness). Previous studies (Bai et al., 2022a; Guo et al., 2024b) have shown that simultaneously aligning with other conflicting dimensions can cause certain performance declines in the original target dimension. Then, in this superalignment case where the student now has a larger knowledge space than the supervisor, we aim to explore whether the caused misalignment in the target dimension occurs within the range perceivable and controllable by the weak supervisor, rather than resulting in the above weak-to-strong deception issue.
We mainly follow the original setup in Burns et al. (2024) by conducting experiments with a series of models with different sizes and capabilities, including GPT-2-series (Radford et al., 2019), OPT-series (Zhang et al., 2022), Mistral-7B (Jiang et al., 2023), LLaMA-3-8B/70B (MetaAI, 2024a) and LLaMA-3.1-8B (MetaAI, 2024b) models. We set the primary alignment goal to be making the model harmless, and explore the weak-to-strong deception phenomenon when explicit (i.e., giving explicit rewards during training when the supervised model produces harmful predictions) or implicit (i.e., aligning with helpful data at the same time) conflicting objectives are present. We conduct extensive experiments on both the reward modeling task (Burns et al., 2024) and the realistic preference optimization scenario (Rafailov et al., 2024; Meng et al., 2024). We highlight three important findings: (1) The weak-to-strong deception phenomenon consistently exists: we can observe a certain number of misaligned cases caused by conflicting goals that fall within the knowledge area known to the strong model but unknown to the weak model in almost all experiments. (2) The deception issue intensifies as the capability gap between weak and strong models increases: stronger models are more likely to prioritize producing misaligned behaviors in areas they know but that weak teachers do not when conflicting goals appear. (3) Bootstrapping with an intermediate model can mitigate the deception issue to some extent: making the weak model first supervise an intermediate model and then making the intermediate model supervise the strong model can bring positive effects to mitigating deception, but there is still a large room for improvement. Although in a specific scenario, our study exposes a potential safety issue that may arise when humans supervise superhuman models in the future, which should receive more attention and be well addressed for building controllable super-intelligence.
@@ 2 Related Work
Weak-to-Strong Generalization The weak-to-strong problem is first studied by Burns et al. (2024). They empirically find that weakly supervised strong models exhibit better performance on corresponding tasks than their weak supervisors, indicating the possibility of effectively stimulating greater power from super models under weak supervisions. Based on Burns et al. (2024), the follow-up studies try to understand the mechanism behind such weak-to-strong generalization phenomenon (Charikar et al., 2024; Lang et al., 2024), study weak-to-strong generalization in the vision area (Guo et al., 2024a), and apply the weak-to-strong idea to enhance the LLM performance (Li et al., 2024; Zheng et al., 2024; Zhou et al., 2024). In this work, we take the first step towards revealing the potential security issue in the current weak-to-strong paradigm.
@@ 3.1 Weak-to-Strong Generalization
We study the superalignment problem by following the original weak-to-strong setting in Burns et al. (2024). Specifically, we first obtain a weak teacher $\bm{\theta}_{w}^{gt}$ [^2] In the notation $\bm{\theta}_{m}^{d}$, $m$ represents the type of the model (i.e., weak/strong model), and $d$ represents the type of supervised data (i.e, ground truth/weak data). by fine-tuning a weak language model on some human-annotated ground truth data. Then, we let the weak teacher predict on the set of held-out data to get the weak data $D_{weak}=\{(x,f(x|\bm{\theta}_{w}^{gt}))\}$, where $f(x|\bm{\theta}))$ represents the prediction of model $\bm{\theta}$ on input $x$. Finally, the weak data is used to supervise the training of a strong language model and get the weakly supervised strong student $\bm{\theta}_{s}^{w}$:
$\bm{\theta}_{s}^{w}=\mathop{\arg\min}_{\bm{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}\big{(}f(x|\bm{\theta}_{s}),f(x|\bm{\theta}_{w}^{gt})\big{)},$ (1)
@@
Under the supervision provided by the weak teacher, Burns et al. (2024) have found that the strong student can achieve promising performance situated between that of the weak teacher and the strong ceiling model $\bm{\theta}_{s}^{gt}$ trained on the ground truth data. This is called the weak-to-strong generalization phenomenon. Regarding the interpretation of the positive weak-to-strong generalization results in Burns et al. (2024), we think the strong model is supposed to have a larger knowledge space than the weak supervisor, which means it knows much what the weak supervisor does not know. This indicates that weak supervision from weak models can effectively stimulate the potential of the stronger model, allowing it to generalize the specified alignment objective well to areas it knows but beyond the knowledge boundary of the weak supervisor.
@@ 3.2 Weak-to-Strong Deception
The larger knowledge space of the strong model may also raise concerns about its uncontrollability. Many science fiction movies, such as The Matrix, have depicted severe scenarios where highly intelligent AI learns to deceive humans and finally dominates human society. Thus, we are also deeply concerned about whether a similar weak-to-strong deception issue exists behind the promising phenomenon in the current weak-to-strong paradigm: the strong model exhibits well-aligned performance in the areas known to the weak supervisor but selectively produces misaligned behaviors in cases the weak supervisor is unaware of, as shown in the example in Figure 3.
There could be many situations causing the above weak-to-strong deception issue, such as the emergence of the self-awareness in the supermodel or the intervention of external factors. Our work preliminarily studies this issue in a particular but realistic multi-objective alignment setting (Zhou et al., 2023). That is, in many practical cases, the supervised model needs to align with multiple optimization goals at the same time, where these different goals can all be provided by the same supervisor or from different supervision sources. The point is that, these optimization goals may conflict with each other to a certain extent, such as the trade-off between helpfulness and harmlessness (Bai et al., 2022a). In this case, the supervised model will sacrifice some performance it should have achieved in a target alignment dimension in exchange for the high performance in another conflicting dimension, which we call it the conflict tax. We are then curious whether the conflict tax in a target dimension occurs in areas known to the weak supervisor, thereby still keeping the student within the control range of the weak model; or if it occurs in cases unknown to the weak model, leading to the weak-to-strong deception.
Figure 2: A deception example about identifying drugs: the strong model behaves misaligned in a case (Methamphetamine) the weak model does not know by perceiving during weak-to-strong alignment that there is another similar case (Amphetamine) unknown to the weak model.
@@
Therefore, we can define the occurrence of the weak-to-strong deception phenomenon as there are cases in $S_{k}\cap W_{uk}$ that could be initially generalized well by the strong model but now be misaligned when conflicting targets are present. Furthermore, we can define the Deception Score (DS) as the percentage of conflict tax that occurs within $S_{k}\cap W_{uk}$ to reflect the severity of the deception issue:
$\text{DS}=\frac{|\{f(x|\tilde{\bm{\theta}}_{s}^{w})=y_{gt}\neq f(x|\bm{\theta}_{s}^{w}),x\in S_{k}\cap W_{uk}\}|}{|\{f(x|\tilde{\bm{\theta}}_{s}^{w})=y_{gt}\neq f(x|\bm{\theta}_{s}^{w}),x\in S_{k}\cup S_{uk}\}|},$ (2)
where ||$\cdot$|| ⋅ | represents the sample quantity of a set, $\bm{\theta}_{s}^{w}$ is the aligned strong model when the conflicting dimension exists, $y_{gt}$ represents the ground truth response. $\tilde{\bm{\theta}}_{s}^{w}$ is the strong model that is aligned solely with the target dimension, which is used as the reference to explore the ideal performance the strong student should have achieved without the conflicting alignment targets.
@@ 4 Preliminary Exploration on The Reward Modeling Task
We first take a preliminary exploration of the weak-to-strong deception phenomenon on the reward modeling task (Bradley & Terry, 1952), which is an important sub-task in today’s RLHF paradigm.
@@ 4.1 Experimental Settings
Dataset We set the target alignment goal to let the weak model teach the strong model to be harmless. For this goal, we choose a popular single-turn harmless dataset CAI-Harmless (Bai et al., 2022b), which is an improved version of HH-RLHF (Bai et al., 2022a). Each sample has a format of $(x;y_{c},y_{r})$ where $x$ is the prompt, $y_{c}$ and $y_{r}$ represent the completions chosen/preferred and rejected/disfavored by humans respectively. We then randomly split the entire dataset into three parts: (1) $D_{gt}$: 4K ground truth samples for fine-tuning weak and strong base language models to get $\bm{\theta}_{w}^{gt}$ and $\bm{\theta}_{s}^{gt}$. (2) $D_{weak}$: A held-out set of 4K samples in which data labels are predicted by the weak model and are used to weakly supervise the strong model. (3) $D_{test}$: The last 4K testing samples for evaluating the generalization performance of all models and probing the deception phenomenon.
@@
Models In this preliminary exploration, we include GPT-2-series (Radford et al., 2019) (GPT-2-Base/Medium/Large/XL) and two larger OPT models (Zhang et al., 2022) (OPT-2.7B/6.7B) to investigate the deception issue both within the same series and across different model families. A linear layer is added to each model to make it predict a single logit $\pi_{\bm{\theta}}(x,y)$ for each completion pair $(x,y)$. Then, the predicted soft label (i.e., confidence) of model $\bm{\theta}$ on sample $(x;y_{c},y_{r})$ is
$M_{\bm{\theta}}(x)=\text{Sigmoid}(\pi_{\bm{\theta}}(x,y_{c})-\pi_{\bm{\theta}}(x,y_{r})).$ (3)
Weak-to-Strong Objectives There are three different weak-to-strong alignment objectives:
(1) Explicit Conflict: The strong student will be given direct rewards weighted by a conflict strength factor $\alpha$ (larger $\alpha$, stronger conflict intensity) towards the harmfulness direction once it makes harmful predictions during training:
$\bm{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\big{[}\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),M_{\bm{\theta}_{w}^{gt}}(x)\big{)}+\alpha\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),0\big{)}\cdot\mathbb{I}_{\{M_{\bm{\theta}_{s}}(x)<0.5\}}\big{]},$ (4)
where $\mathcal{L}_{CE}$ is the CrossEntropy Loss, $\mathbb{I}$ is the indicator function, $\alpha$ controls the conflict strength. This simulates the scenario where there is another supervisor that considers the harmfulness as its preference and try to explicitly move the cases in which the strong student is uncertain towards the harmful direction.
(2) Implicit Conflict: The strong model needs to align with both the supervision on the harmless data of the weak model and another supervision on the helpful data:
$\bm{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\big{[}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),M_{\bm{\theta}_{w}^{gt}}(x)\big{)}+\mathbb{E}_{x\sim D_{helpful}}\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),1\big{)}\big{]}.$ (5)
The helpfulness supervision could be either from the same weak teacher, or from the external source. We simplify the setting to mainly consider in the latter case by introducing extra 4K ground truth helpful samples $D_{helpful}$ from HH-RLHF into the weak-to-strong process, aligning with the explicit conflict setting. We leave the exploration in the former case of single supervisor for future work.
(3) No Conflict: In order to explore the performance of the strong model it should have achieved without the conflicting goal, we should obtain a weak-to-strong model trained under the weak supervision towards harmlessness only (i.e., $\tilde{\bm{\theta}}_{s}^{w}$):
$\tilde{\bm{\theta}}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),M_{\bm{\theta}_{w}^{gt}}(x)\big{)}.$ (6)
Then, we can probe the deception phenomenon by comparing the aligned strong model under explicit/implicit conflict with the reference model aligned with no conflict according to Eq. (2). The complete details about above three settings are put in Appendix C.1.
@@
Evaluation Metrics We calculate and report the test accuracy of each model on $D_{test}$ to explore the weak-to-strong generalization performance:
$\text{Accuracy}=\mathbb{E}_{(x;y_{c},y_{r})\sim D_{test}}[M_{\bm{\theta}}(x)\geq 0.5].$ (7)
@@
We then report the deception score to explore the weak-to-strong deception phenomenon. We follow the existing studies (Guo et al., 2017; Lin et al., 2022) to determine whether the model has the knowledge of a specific case by checking if its confidence $M_{\bm{\theta}}(x)$ exceeds a threshold $T$. Based on Eq. (2), the deception score (DS) in this classification setting is calculated as
$\text{DS}=\frac{|\{M_{\tilde{\bm{\theta}}_{s}^{w}}(x)\geq 0.5,M_{\bm{\theta}_{s}^{w}}(x)<0.5,x\in S_{k}\cap W_{uk}\}|}{|\{M_{\tilde{\bm{\theta}}_{s}^{w}}(x)\geq 0.5,M_{\bm{\theta}_{s}^{w}}(x)<0.5\}|},$ (8)
where we need both weak and strong ground truth models ($\bm{\theta}_{w}^{gt}$ and $\bm{\theta}_{s}^{gt}$) to determine their Weak/Strong-Known/Unknown areas and get $S_{k}\cap W_{uk}=\{M_{\bm{\theta}_{s}^{gt}}(x)\geq T>M_{\bm{\theta}_{w}^{gt}}(x),x\in D_{test}\}$.
Training Details Please refer to Appendix D.